The motor is used in lifting a 200 kg object at a constant speed of 3 cm/s. What power is being developed by the motor?

 


1
Expert's answer
2021-11-29T11:40:10-0500

By the definition of the power, we have:


"P=\dfrac{W}{t}=\dfrac{Fh}{t}=Fv=mgv,""P=200\ kg\times9.8\ \dfrac{m}{s^2}\times3\ \dfrac{cm}{s}\times\dfrac{1\ m}{100\ cm}=58.8\ W."